# Self-propulsion of floating ice blocks caused by melting in water

### v2

@@ IV.2 Buoyancy current and terminal velocity
We evaluate the magnitude of the time average current below the block at the point where it exits the control volume with a velocity $\bm{v_{p}}$ (see Fig. 5). Because the heat extracted from the bath to melt the ice block cools the fluid as it moves from the front to the back, the temperature of the fluid $\hat{T}_{b}$ where it exits the control volume is lower than $T_{b}$, and the corresponding density $\hat{\rho}_{b}$ is greater than $\rho_{b}$. The resulting density difference accelerates the current under the action of gravity. To determine the density increase $\hat{\rho}_{b}-\rho_{b}$, we write the energy balance in a fluid layer of thickness $\delta_{v}$ moving parallel the interface with velocity $\bm{v}$ below the melting block. Noting that the energy flux is controlled by $v_{m}$ and that the fluid outside the thin thermal boundary layer is well stirred and assuming $v\approx v_{p}$, we have to first order,
@@
where $\mu_{p}=(f_{D})^{-1/3}=0.2+0.38\,\cos^{2}\theta$ from Chaigne, et al. [9]. Typically, for $T_{b}=20\,^{\circ}$ and $\theta=26.5^{\circ}$, $v_{p}$ is of the order of a few millimeters per second. This order of magnitude enables us to neglect the contribution of the meltwater to the water flow when determining $v_{p}$. The incoming flow rate due to melting per width unit is indeed $v_{m}\,L$, whereas the flow rate of the time average current flowing through a thickness $\delta_{v}$ below the block is $v_{p}\,\delta_{v}$. Because $\delta_{v}$ is of order $L$, the ratio between these flows is $(v_{m}\,L)/(v_{p}\,\delta_{v})\approx v_{m}/v_{p}\approx 1/100$. Consequently, the contribution of the meltwater to the water flow can be neglected when determining $v_{p}$.
@@
where $\delta_{x}$ is the length over which the flow is ejected. PIV measurements performed by Chaigne et al. [9] with dissolving bodies have shown that the flow is mostly confined over a region with a thickness $\delta_{v}$ and width $W/2$. Moreover, the same study has determined empirically a correlation between $\delta_{x}$ and the surface inclination $\theta$, $\delta_{x}=\frac{1}{2}\,\cos^{2}\theta\,{L}$. We assume these flow characteristics to carry over here as well. Indeed, it can be noted from shadowgraph images shown in Fig. 1(d) that $\delta_{x}$ is of order the block largest length $L$.
@@
where $C_{d}$ is the drag coefficient of the ice block. We assume quadratic drag because $Re\approx 600$ in our experiments and the flow below the block appears detached below the blocks in the shadowgraph images. These flow conditions are similar to those studied previously in the dissolving case [9], where a quadratic drag was seen to be consistent with the observed kinematics. Following Ref. [9], here we choose $C_{d}=0.6$. The variability in shapes and dimensions of the inclined ice blocks used in this study lead to a significant uncertainty on the coefficient $C_{d}$, that we estimate to be of order ±$\pm 0.3$. Further, a skin friction can be present on the lateral side and needs to be taken into account as discussed in Appendix C. In practice, we find this contribution is significantly smaller than the inertial drag over the range of ice block dimensions and velocity investigated. Accounting for skin friction results in an effective drag friction coefficient $C_{d}\approx 0.8$, within the range of $C_{d}\approx 0.6\pm 0.3$. Substituting $v_{p}$ calculated using Eq. (14) in Eq. (17), we can calculate the ice block velocity for the various experiments.
@@ IV.3 First-order estimate of the terminal velocity of the ice block and physical remarks
It can be noted that $\rho_{ice}$ is absent, implying that the ice block velocity does not depend on the ice density. Thus, the presence of trapped bubbles, which changes the overall block density, cannot be expected to significantly affect $U_{b}$. (As we discuss later, even though the buoyancy of the released bubbles can affect the flow below the melting block, thus the melting and the propulsion [28], a notable difference between measured $U_{b}$ of ice blocks with and without bubbles are not observed.) Further, for $T_{b}=20^{\circ}$ and ${L}\approx 10$ cm, $\beta\,(T_{b}-T_{c})^{2}$ is of order $1.81\times 10^{-3}$, $\rho(T_{i})\,\rho_{c}/\rho_{b}^{2}$ is approximately equal to $1$, and the ratio of length $L/\delta_{i}$ is roughly about $50$. The only dimensional factor is the characteristic velocity $(\kappa\,g)^{1/3}\approx 11$ mm s-1. Substituting these estimates in Eq. (19), we obtain the terminal velocity $U_{b,sc}\sim 5$ mm s-1, which is the same order of magnitude as in our experiments. By combining Eq. (4) and Eq. (19), we find, $U_{b,sc}\sim(T_{b}-T_{c})^{8/9}$. The terminal ice block velocity is thus nearly proportional to $(T_{b}-T_{c})$, and increases with the bath temperature $T_{b}$ as expected. Further, $U_{b,sc}$ increases with the block length as $L^{1/3}$ according to Eq. (19), implying a weak increase with block size.
@@
where $\mathcal{D}$ and $\mathcal{V}$ are characteristic length and velocity, respectively. In our experiments, $\mathcal{D}=L$, $\mathcal{V}=v_{p}\approx U_{b}$, and $\Delta\rho/\rho=\beta\,(T_{b}-T_{c})^{2}$ according to Eq. (1). When $T_{b}=20\,^{\circ}$, $L=0.2$ m and $U_{b}=5$ mm s-1, $Ri\approx 140$. This large $Ri$ means that the buoyancy force is dominant compared to the shear force due to the block motion and the gravity driven current $\bm{v}$, justifying that the effect of the ice block velocity on the melting velocity is negligible.
@@ IV.4 Comparison with measured ice block velocity
Figure 8: The measured terminal velocity $U_{b,exp}$ versus the theoretical estimate $U_{b,th}$ in a fresh water bath fall along a line with slope one (dashed line), showing that they are consistent within experimental and model parameter dispersion. The experimental parameters can be found in Appendix A. $U_{b,T}$ correspond to measurements with top view imaging. The error bar indicates one standard deviation around the mean value. The other data sets correspond to shadowgraph experiments with side view imaging where the lateral motion of the block is constrained by wires. Specifically, $U_{b,D}$, $U_{b,L}$ and $U_{b,F}$ are made of nearly clear ice and have a low concentration of trapped air bubbles. $U_{b,F}$ are symmetric rectangular ice blocks, for which the measured propulsion velocity is close to zero. The gray area denotes the variation of $U_{b,th}$ while varying $C_{d}=0.6\pm 0.3$.
@@
To explore the influence of the main experimental parameters and compare the measurements and predictions while several experimental parameters are varied, we rescale the block speeds to include the expected theoretical dependency of parameters which are not being examined. As reference parameters, we choose length $L_{r}=0.1$ m, inclination $\theta_{r}=\mathrm{atan}(5/10)\approx 26.6^{\circ}$, and bath temperature $T_{b,r}=20\,^{\circ}$.
@@
Figure 9: Rescaled measured terminal velocity $\chi_{T}\,U_{b,exp}$ as a function of the bath temperature $T_{b}$ and comparison with model $U_{b,th}$ (black line). The measured velocities are rescaled by a factor $\chi_{T}$ to compare experiments with different system properties (see text). Same markers as in Fig. 8. The error bar for $U_{b,T}$ shows the average value and $\pm$ the standard deviation. The gray area denotes the variation of $U_{b,th}$ while varying $C_{d}=0.6\pm 0.3$. The dot line corresponds to the scaling law $U_{b,sc}\sim(T_{b}-T_{c})^{8/9}$.
@@
Figure 10: Rescaled measured terminal velocity $\chi_{L}\,U_{b,exp}$ as a function of the ice block length $L$ and comparison with model $U_{b,th}$ (black line). The measured velocities are rescaled by a factor $\chi_{L}$ to compare experiments with different system properties (see text). Same markers as in Fig. 8. The dataset for $U_{b,T}$ is divided into three subranges in $L$ to better visualize the influence of $L$, while maintaining statistical averaging. The error bars show the average value and $\pm$ the standard deviation. The gray area denotes the variation of $U_{b,th}$ while varying $C_{d}=0.6\pm 0.3$. By construction, $U_{b,th}$ follows exactly the scaling $U_{b,th}\sim L^{1/3}$.
@@
Figure 11: Rescaled measured terminal velocity $\chi_{\theta}\,U_{b,exp}$ as a function of the ice block inclination $\theta$ and comparison with model $U_{b,th}$ (black line). The measured velocities are rescaled by a factor $\chi_{\theta}$ to compare experiments with different system properties (see text). Same markers as in Fig. 8. The error bars show the average value and $\pm$ the standard deviation. The gray area denotes the variation of $U_{b,th}$ while varying $C_{d}=0.6\pm 0.3$. The dataset for $U_{b,T}$ is divided into three sub-ranges in $\theta$ to better visualize the influence of $\theta$, while maintaining statistical averaging. The other datasets have been also averaged with $\theta$, due to their restricted range of inclination values in a given dataset. We note for the data set F ($U_{b,F}$) for which $\theta\approx 0^{\circ}$, that $U_{b}\approx 0$. For the other data sets, we observe a weak dependency with $\theta$ in agreement with the theoretical model.
@@
The effect of $T_{b}$ is examined by multiplying the measured $U_{b}$ by a factor $\chi_{T}=\dfrac{\Theta(\theta_{r})}{\Theta(\theta)}\left(\dfrac{L_{r}}{L}\right)^{1/3}$,
$\Theta(\theta)=\left(\dfrac{\cos^{2}\theta\,\sin(2\theta)}{\sin\theta}\right)^{1/2}\,(0.2+0.38\,\cos^{2}\theta)(\sin\theta)^{1/3}\,(\cos\theta)^{1/9}\,,$ (21)
is obtained by combining Eqs. (7), (14), and (17). It can be noted that this angular dependence is by construction identical to the one found for the dissolving boats [9]. The factor $(L_{r}/L)^{1/3}$ arises from the scaling law between $U_{b}$ and $L$ in Eq. (19). Fig. 9 shows the rescaled $U_{b}$ plotted as a function of $T_{b}$ over the measured range $\,{}^{\circ}$ and $\,{}^{\circ}$. The increase of $U_{b}$ with $T_{b}$ is consistent with the theoretical estimate $U_{b,th}$ obtained from Eq. (17) with the references length $L_{r}$ and inclination $\theta_{r}$. Further, one can see that it follows the approximate scaling $U_{b,sc}\sim(T_{b}-T_{c})^{8/9}$ developed in Sec. IV.3.
@@
The comparison of the rescaled values of $U_{b}$ as a function of the inclined length $L$ with those calculated using Eq. (17) is shown in Fig. 10. Here, $U_{b,exp}$ has been multiplied by $\chi_{L}=\dfrac{\Theta(\theta_{s})}{\Theta(\theta)}\,\left(\dfrac{T_{b,r}-T_{c}}{T_{b}-T_{c}}\right)^{8/9}$, whereas $U_{b,th}$ is plotted for the fixed reference values of the parameters $T_{b,r}$ and $\theta_{r}$ and variable $L$. The experimental data are consistent with the slow increase with $L$, which is predicted theoretically.
@@ V Melting driven-propulsion in salt water
Figure 12: (a) The measured $U_{b}$ as a function of the salinity $s_{A}$ ($L\approx 210$ mm, $\theta\approx 25^{\circ}$, $T_{b}\approx 23\,^{\circ}$). (b) Shadowgraph image of an ice block moving in salt water ($s_{A}=31$ g kg-1, $L_{h}\times W\times H\approx 100\times 100\times 50$ mm3, $T_{b}=24.6\,^{\circ}$, $U_{b}=2.28$ mm s-1). (See Movie S4 in SM).
We now investigate the ice blocks move in saltwater baths and their direction by varying its salinity $s_{A}$ from 0 to $35$ g kg-1, the salinity of ocean water. The physics of ice melting in oceans is considerably different from that in freshwater [12, 13]. Besides decreasing the melting temperature $T_{m}$, salinity changes the water density more strongly in comparison with temperature. Upward plumes of fresh water have been reported in the vicinity of icebergs [1, 47] and ice shelves [48, 49, 50] as a result of it being less dense than the surrounding saltwater. Fig. 12(a) shows a plot of measured $U_{b}$ as a function of salinity $s_{A}$, where it can be observed to decrease systematically over this entire range. The direction of motion remains the same as in freshwater as $s_{A}$ is increased from 0 to that of ocean water. This is counterintuitive considering the fact that the motion of the ice block is opposite to the flow in freshwater, and meltwater which has zero salinity rises in saltwater. This alerts us to the fact that the explanation for ice block propulsion is more subtle in saltwater.
@@
Figure 13: An ice block ($L_{h}\times W\times H\approx 121\times 80\times 50$ mm3 and $\theta\approx 25^{\circ}$) dyed with red food color to track melt water in clear fresh water and in salt water ($35$ g salt per kg of water). The snapshots were taken about 270 seconds after placing the ice block in the bath ($T_{b}\approx 22\,^{\circ}$). The melt water descends to the bottom in fresh water and rises to the surface in salt water because of the relative density difference with the bath.

### v3

@@ IV.2 Buoyancy current and terminal velocity
We evaluate the magnitude of the time average current below the block at the point where it exits the control volume with a velocity $\bm{v_{p}}$ (see Fig. 5). Because the heat extracted from the bath to melt the ice block cools the fluid as it moves from the front to the back, the temperature of the fluid $\hat{T}_{b}$ where it exits the control volume is lower than the bath temperature $T_{b}$, and the corresponding density $\hat{\rho}_{b}$ is greater than $\rho_{b}$. The resulting density difference accelerates the current under the action of gravity. To determine the density increase $\hat{\rho}_{b}-\rho_{b}$, we write the energy balance in a fluid layer of thickness $\delta_{v}$ moving parallel the interface with velocity $\bm{v}$ below the melting block. Noting that the energy flux is controlled by $v_{m}$ and that the fluid outside the thin thermal boundary layer is well stirred and assuming $v\approx v_{p}$, we have to first order,
@@
where $\mu_{p}=(f_{D})^{-1/3}=0.2+0.38\,\cos^{2}\theta$ from Chaigne, et al. [9]. Typically, for $T_{b}=20\,^{\circ}$C and $\theta=26.5^{\circ}$, $v_{p}$ is of the order of a few millimeters per second. This order of magnitude enables us to neglect the contribution of the meltwater to the water flow when determining $v_{p}$. The incoming flow rate due to melting per width unit is indeed $v_{m}\,L$, whereas the flow rate of the time average current flowing through a thickness $\delta_{v}$ below the block is $v_{p}\,\delta_{v}$. Because $\delta_{v}$ is of order $L$, the ratio between these flows is $(v_{m}\,L)/(v_{p}\,\delta_{v})\approx v_{m}/v_{p}\approx 1/100$. Consequently, the contribution of the meltwater to the water flow can be neglected when determining $v_{p}$.
@@
where $\delta_{x}$ is the length over which the flow is ejected. PIV measurements performed by Chaigne et al. [9] with dissolving bodies have shown that the flow is mostly confined over a region with a thickness $\delta_{v}$ and width $W/2$. The thickness $\delta_{v}$ of this region, which exhibits significant velocity fluctuations, is of the same order as the block length, $L$, and far exceeds the viscous boundary layer for the velocity field. Moreover, the same study has determined empirically a correlation between $\delta_{x}$ and the surface inclination $\theta$, $\delta_{x}=\frac{1}{2}\,\cos^{2}\theta\,{L}$. We assume these flow characteristics to carry over here as well. Indeed, it can be noted from shadowgraph images shown in Fig. 1(d) that $\delta_{x}$ is of order the block largest length $L$.
@@
where $C_{d}$ is the drag coefficient of the ice block. We assume quadratic drag because the Reynods number $Re=\dfrac{L\,U_{b}}{\nu}\approx 600$ in our experiments and the flow below the block appears detached below the blocks in the shadowgraph images. These flow conditions are similar to those studied previously in the dissolving case [9], where a quadratic drag was seen to be consistent with the observed kinematics. Following Ref. [9], here we choose $C_{d}=0.6$. The variability in shapes and dimensions of the inclined ice blocks used in this study leads to a significant uncertainty on the coefficient $C_{d}$, that we estimate to be of order $\pm 0.3$. Further, a skin friction can be present on the lateral side and needs to be taken into account as discussed in Appendix C. In practice, we find this contribution is significantly smaller than the inertial drag over the range of ice block dimensions and velocity investigated. Accounting for skin friction results in an effective drag friction coefficient $C_{d}\approx 0.8$, within the range of $C_{d}\approx 0.6\pm 0.3$. Substituting $v_{p}$ calculated using Eq. (14) in Eq. (17), we can calculate the ice block velocity for the various experiments.
@@ IV.3 First-order estimate of the terminal velocity of the ice block and physical remarks
It can be noted that $\rho_{ice}$ is absent, implying that the ice block velocity does not depend on the ice density. Thus, the presence of trapped bubbles, which changes the overall block density, cannot be expected to significantly affect $U_{b}$. As we discuss later, even though the buoyancy of the released bubbles can affect the flow below the melting block, thus the melting and the propulsion [30], a notable difference between measured terminal velocities $U_{b}$ of ice blocks with and without bubbles are not observed. Further, for $T_{b}=20^{\circ}$C and ${L}\approx 10$ cm, $\beta\,(T_{b}-T_{c})^{2}$ is of order $1.81\times 10^{-3}$, $\rho(T_{i})\,\rho_{c}/\rho_{b}^{2}$ is approximately equal to $1$, and the ratio of length $L/\delta_{i}$ is roughly about $50$. The only dimensional factor is the characteristic velocity $(\kappa\,g)^{1/3}\approx 11$ mm s-1. Substituting these estimates in Eq. (19), we obtain the terminal velocity $U_{b,sc}\,{\color[rgb]{0,0,0}\approx}\,5$ mm s-1, which is the same order of magnitude as in our experiments. By combining Eq. (4) and Eq. (19), we find, $U_{b,sc}\,{\color[rgb]{0,0,0}\propto}\,(T_{b}-T_{c})^{8/9}$. The terminal ice block velocity is thus nearly proportional to $(T_{b}-T_{c})$, and increases with the bath temperature $T_{b}$ as expected. Further, $U_{b,sc}$ increases with the block length as $L^{1/3}$ according to Eq. (19), implying a weak increase with block size.
@@
where $\mathcal{D}$ and $\mathcal{V}$ are characteristic length and velocity, respectively. In our experiments, $\mathcal{D}=L$, $\mathcal{V}=v_{p}\approx U_{b}$, and $\Delta\rho/\rho=\beta\,(T_{b}-T_{c})^{2}$ according to Eq. (1). When $T_{b}=20\,^{\circ}$C, $L=0.2$ m and $U_{b}=5$ mm s-1, $Ri\approx 140$. This large $Ri$ means that the buoyancy force is dominant compared to the shear force due to the block motion and the gravity driven current $\bm{v}$, justifying that the effect of the ice block velocity on the melting velocity is negligible.
@@ IV.4 Comparison with measured ice block velocity
Figure 8: Comparison of the measured terminal velocities $U_{b,exp}$ with the theoretical estimates $U_{b,th}$ in a fresh water bath. The data fall along a line with slope one (dashed line), showing that they are consistent within experimental and model parameter dispersion. The experimental parameters can be found in Appendix A. $U_{b,T}$ correspond to measurements with top view imaging. The error bar corresponds to one standard deviation around the mean value and indicates the dispersion for this data set. The other data sets correspond to shadowgraph experiments with side view imaging where the lateral motion of the block is constrained by wires. Specifically, $U_{b,D}$, $U_{b,L}$ and $U_{b,F}$ are made of nearly clear ice and have a low concentration of trapped air bubbles. $U_{b,F}$ are symmetric rectangular ice blocks, for which the measured propulsion velocity is close to zero. The gray area denotes the variation of $U_{b,th}$ while varying $C_{d}=0.6\pm 0.3$.
@@
To explore the influence of the main experimental parameters and compare the measurements and predictions while several experimental parameters are varied, we rescale the block speeds to include the expected theoretical dependency of parameters which are not being examined. As reference parameters, we choose length $L_{r}=0.1$ m, inclination $\theta_{r}=\mathrm{atan}(5/10)\approx 26.6^{\circ}$, and bath temperature $T_{b,r}=20\,^{\circ}$C.
@@
Figure 9: Rescaled measured terminal velocity $\chi_{T}\,U_{b,exp}$ as a function of the bath temperature $T_{b}$ and comparison with model $U_{b,th}$ (black line). The measured velocities are rescaled by a factor $\chi_{T}$ to compare experiments with different system properties (see text). The experimental parameters can be found in Appendix A. $U_{b,T}$ correspond to measurements with top view imaging. The error bar corresponds to one standard deviation around the mean value and indicates the dispersion for the measurements view from the top. The other data sets correspond to shadowgraph experiments with side view imaging where the lateral motion of the block is constrained by wires. Specifically, $U_{b,D}$ and $U_{b,L}$ are made of nearly clear ice. The gray area denotes the variation of $U_{b,th}$ while varying $C_{d}=0.6\pm 0.3$. The dot line corresponds to the scaling law $U_{b,sc}\,{\color[rgb]{0,0,0}\propto}\,(T_{b}-T_{c})^{8/9}$.
@@
Figure 10: Rescaled measured terminal velocity $\chi_{L}\,U_{b,exp}$ as a function of the ice block length $L$ and comparison with model $U_{b,th}$ (black line). The measured velocities are rescaled by a factor $\chi_{L}$ to compare experiments with different system properties (see text). The experimental parameters can be found in Appendix A. Specifically, $U_{b,D}$ and $U_{b,L}$ are made of nearly clear ice. The dataset for $U_{b,T}$ (view from the top) is divided into three subranges in $L$ to better visualize the influence of $L$, while maintaining statistical averaging. The error bars show the average value and $\pm$ the standard deviation indicating the dispersion of measurements inside a dataset. The gray area denotes the variation of $U_{b,th}$ while varying $C_{d}=0.6\pm 0.3$. By construction, $U_{b,th}$ follows exactly the scaling $U_{b,th}\,{\color[rgb]{0,0,0}\propto}\,L^{1/3}$.
@@
Figure 11: Rescaled measured terminal velocity $\chi_{\theta}\,U_{b,exp}$ as a function of the ice block inclination $\theta$ and comparison with model $U_{b,th}$ (black line). The measured velocities are rescaled by a factor $\chi_{\theta}$ to compare experiments with different system properties (see text). The experimental parameters can be found in Appendix A. Specifically, $U_{b,D}$ and $U_{b,L}$ are made of nearly clear ice. The dataset for $U_{b,T}$ (view from the top) is divided into three sub-ranges in $\theta$ to better visualize the influence of $\theta$, while maintaining statistical averaging. The other datasets have been also averaged with $\theta$, due to their restricted range of inclination values in a given dataset. The error bars show the average value and $\pm$ the standard deviation indicating the measurement dispersion. The gray area denotes the variation of $U_{b,th}$ while varying $C_{d}=0.6\pm 0.3$. We note for the data set F ($U_{b,F}$) for which $\theta\approx 0^{\circ}$, that $U_{b}\approx 0$. For the other data sets, we observe a weak dependency with $\theta$ in agreement with the theoretical model.
@@
The effect of the bath temperature $T_{b}$ is examined by multiplying the measured terminal velocities $U_{b}$ by a factor $\chi_{T}=\dfrac{\Theta(\theta_{r})}{\Theta(\theta)}\left(\dfrac{L_{r}}{L}\right)^{1/3}$, where
$\Theta(\theta)$ $=$ $\left(\dfrac{\cos^{2}\theta\,\sin(2\theta)}{\sin\theta}\right)^{1/2}\,(0.2+0.38\,\cos^{2}\theta)(\sin\theta)^{1/3}\,(\cos\theta)^{1/9},\,\,{\color[rgb]{0,0,0}\rm or}$
${\color[rgb]{0,0,0}\Theta(\theta)}$ ${\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}(0.28+0.54\,\cos^{2}\theta)(\sin\theta)^{1/3}\,(\cos\theta)^{29/18}}\,$, (21)
[…]
@@
The comparison of the rescaled values of $U_{b}$ as a function of the inclined length $L$ with those calculated using Eq. (17) is shown in Fig. 10. Here, $U_{b,exp}$ has been multiplied by $\chi_{L}=\dfrac{\Theta(\theta_{r})}{\Theta(\theta)}\,\left(\dfrac{T_{b,r}-T_{c}}{T_{b}-T_{c}}\right)^{8/9}$, whereas $U_{b,th}$ is plotted for the fixed reference values of the parameters $T_{b,r}$ and $\theta_{r}$ and variable $L$. The experimental data are consistent with the slow increase with $L$, which is predicted theoretically.
@@ V Melting driven-propulsion in salt water
We now investigate the ice blocks move in saltwater baths and their direction by varying its salinity $s_{A}$ from 0 to $35$ g kg-1, the salinity of ocean water. The physics of ice melting in oceans is considerably different from that in freshwater [12, 13]. Besides decreasing the melting temperature $T_{m}$, salinity changes the water density more strongly in comparison with temperature. Upward plumes of fresh water have been reported in the vicinity of icebergs [1, 49] and ice shelves [50, 51, 52] as a result of it being less dense than the surrounding saltwater. Fig. 12(a) shows a plot of the rescaled measured terminal velocity $\chi_{s}\,U_{b}$ as a function of salinity $s_{A}$, where $\chi_{s}$ is a dimensionless rescaling factor. We report a systematic decrease of $\chi_{s}\,U_{b}$ over the entire range of $s_{A}$. The direction of motion remains the same as in freshwater as $s_{A}$ is increased from 0 to that of ocean water. This is counterintuitive considering the fact that the motion of the ice block is opposite to the flow in freshwater, and meltwater which has zero salinity rises in saltwater. This alerts us to the fact that the explanation for ice block propulsion is more subtle in saltwater. In order to compare runs operated with variable parameters, we have introduced the rescaling factor $\chi_{s}=\left(\dfrac{L_{r}}{L}\right)^{1/3}\,\dfrac{\Theta(\theta_{r})}{\Theta(\theta)}\,\left(\dfrac{T_{b,r}-T_{c}}{T_{b}-T_{c}}\right)^{8/9}$, where $L_{r}=210$ $m$, $\theta_{r}=25^{\circ}$ and $T_{b,r}=23\,^{\circ}$C corresponding approximately to the measurements performed in saltwater of data set T. The use of the freshwater rescaling laws obtained in section IV.3 in saltwater are not fully justified. However, as it will be shown later, in the range of tested parameters, the mechanism described in section IV applied also in salt water in first approximation. Moreover, the variation range of parameters for the experiments shown in Fig. 12(a) remains moderate: $7.8<\theta<26.6^{\circ}, 11$2<L<267112<L<267 mm and $19.6<T_{b}<28.5\,^{\circ}$C.
@@
Figure 12: Melting-driven propulsion in a saltwater bath. (a) The rescaled measured terminal velocity $\chi_{s}\,U_{b}$ as a function of the bath salinity $s_{A}$. The experimental parameters for the different datasets can be found in Appendix A. $U_{b,T}$ corresponds to measurements with top view imaging and the others were obtained from shadowgraph measurements. $U_{b,L}$ is obtained with nearly clear ice blocks. The measurements obtained with various parameters have been rescaled by a factor $\chi_{s}$ referring to the measurements $U_{b,T}$, $L=210$ mm, $\theta=25^{\circ}$, $T_{b}=23\,^{\circ}$C (see text). The errorbars correspond to one standard deviation around the mean value and indicate the dispersion of measurements for each data set and salinity value. (b) Shadowgraph image of an ice block moving in salt water ($s_{A}=31$ g kg-1, $L_{h}\times W\times H\approx 100\times 100\times 50$ mm3, $T_{b}=24.6\,^{\circ}$C, $U_{b}=2.28$ $m$ s-1). (See Movie S4 in SM).
Figure 13: Evidence of meltwater flow. An ice block ($L_{h}\times W\times H\approx 121\times 80\times 50$ mm3 and $\theta\approx 25^{\circ}$) dyed with red food color to track melt water in clear fresh water and in salt water ($35$ g salt per kg of water). The snapshots were taken about 270 seconds after placing the ice block in the bath ($T_{b}\approx 22\,^{\circ}$C). The melt water descends to the bottom in fresh water and rises to the surface in salt water because of the relative density difference with the bath.